# A detailed dive into fitting strategies for GRB afterglows with contamination: A case study with kilonovae

### v1

@@ 4.1 Scenario 1 - On-axis afterglow with significant kilonova contamination
As discussed in Sec. 3, an on-axis GRB system is likely to suffer from significant kilonova contamination. As the cases of GRB211211A and GRB230307A already prove, they are also the most observationally relevant. We first focus on the strategy where the kilonova is fit independently to the data. The fitted lightcurve from this strategy is shown in Fig. 4 (panel B), which highlights the poor recovery relative to the true input. However, we emphasize that in real observation scenarios, true parameter values are not known, so one could not directly assess the accuracy of inference in this manner. If we disregard the true input curves, the posteriors in Fig. 4 appear to reasonably match the relevant data points for the transient being fit, suggesting a ‘good’ fit to the data that may lead to the conclusion that the inferred parameters are correct.
@@
In Fig. 5, we show violin plots for relevant kilonova parameters from a fit to the data with just a kilonova model in orange, with injection values marked by . Despite producing good fits by eye to the data, our inferred parameters are significantly biased. In particular, and the ejecta mass where the bias is and , respectively (relative to the statistical errors), with no posterior overlap with the true input. Given that these parameters are the most important for interpreting a kilonova, the error seen here suggests that solely fitting the kilonova is not a good strategy.
@@
As shown in Fig. 4 (panel A), the independent fit to the afterglow is a more viable strategy in producing a better match to the true input. This is largely a product of the fit, including the X-ray and radio data, which provide some “calibration” for the afterglow model as the kilonova model has flux at these wavelengths. This motivates the use of the commonly-used strategy of attempting to isolate the kilonova signal by first subtracting the predicted afterglow, and then performing a fit. We approach this “subtraction” strategy by simply masking the X-ray and radio data points and then subtracting the afterglow contribution with the best-fit model parameters for all other wavelengths. To account for the uncertainty in the subtraction, we associate with each subtracted data point the same error as the original data point.
@@
In Fig. 5, we show the posteriors obtained by fitting the kilonova model to the data, subtracting the best-fit afterglow model (in orange). While these posteriors are somewhat more consistent with the input than the independent, kilonova-only fit (cyan), they are still significantly biased. Notably, again, neither the kilonova ejecta mass nor shell velocity is consistent with the input, highlighting the pitfalls of this commonly used strategy. We note that increasing the size of the (largely arbitrary) uncertainty on the error bars when fitting the afterglow-subtracted data could make the kilonova parameters more consistent with the input. However, from a Bayesian perspective, the “subtraction” approach is incorrect and an example of a misspecified model. In particular, when we first fit the afterglow model to data, to maximise the likelihood, the afterglow will depart from the ‘truth’, especially when radio and X-ray data errors are high or the quality otherwise due to cadence or spectral coverage is low relative to the optical. This biased afterglow recovery is then built into the kilonova parameter estimation. Given the intrinsic kilonova and afterglow models are themselves uncertain (unlike our simulation where we know the input), the subtraction strategy can cause significant issues with interpretations. By contrast, the proper (jointly inferring) Bayesian approach to this problem produces kilonova and afterglow posteriors that are marginalising over the other (afterglow) model as opposed to being conditioned on it, and so they will naturally be consistent with the input (given adequate sampling). Moreover, by marginalising over the other model parameters, it is far more likely to yield robust estimated kilonova parameters, as opposed to the subtraction approach.
@@
To verify that the joint approach works, we show fits to the data in Fig. 4 (panel C), which shows that the independent kilonova and afterglow models are consistent with the input. A full corner plot for the joint recovery can be found in Fig. 10 of the appendix, which confirms an accurate recovery with all true kilonova and afterglow parameters recovered correctly to within 3 of the predicted values. We also emphasize that the joint-fitting approach is not significantly more computationally intensive than independently performing an afterglow-only and, subsequently, kilonova-only fit. And importantly, it does not require extra data processing to “generate” afterglow-subtracted kilonova data.
@@ 4.2 Scenario 2 - Off-axis afterglow with significant kilonova contamination
Following the same analysis approaches as Sec. 4.1, we again compare individual, including subtracted, and joint fits for an off-axis GRB system. In Sec. 3 we noted that at early observation epochs, the off-axis afterglow flux density is typically low, which allows the kilonova to dominate the signal in all wavelengths. This suggests that strategies which involve individually fitting the kilonova to the data may be more suited towards the off-axis contamination scenario than the on-axis scenario. The result of our individually fit kilonova component relative to the true inputs, shown in Fig. 4 (panel E), seems to reflect this.
@@
In the lower panel of Fig. 6, we show a violin plot in blue for the individual kilonova recovery. The biases on ejecta mass and shell velocity are reduced in comparison to the on-axis GRB case, discussed in Sec. 4.1. However, the posterior of remains inconsistent with our input. We additionally see that opacity is incorrectly inferred, which is likely due to sensitivity to afterglow contamination around the knee of the lightcurve, which is typically the source of constraints on opacity (alongside the peak time) for the kilonova models used in this work. The inaccuracy of the recovered posteriors despite a seemingly good fit to relevant data points, highlights the need for caution in interpreting fits to subtracted data, and general caution for detailed parameter estimates of quantities with this fitting strategy.
@@
We now look at the individual afterglow fit to data and note that while the kilonova is largely dominant in the lower frequency bandpasses, up to 10 days in lsst-r, this does not necessarily detriment the afterglow recovery. This is again due to the inclusion of uncontaminated radio and X-ray observations which provide a level of calibration to the fit. In the upper panel of Fig. 6, we show violin plots for the key afterglow parameters which we identified as significantly influential to contamination in Sec. 3. Although these properties are recovered at a 3 level (relative to statistical errors), the full recovery shows that there are greater biases on other parameters such as extinction and initial lorentz factor, which are largely informed by the optical/NIR light curve. These inconsistencies again, prompt a review of methods to obtain more accurate predictions so that one may better interpret the afterglow.
@@
If we choose to subtract the best fit afterglow from the data to isolate the kilonova, the resulting inference still fails to correctly recover the input value of ejecta mass. Instead, we may again turn to the use of a combined model. We show the joint fit in panel F of Fig. 4, and note that it provides a good visual fit to the true curves of both afterglow and kilonova components. The full simultaneous inference results show that every parameter is not only recovered correctly, but the predictions have also improved in accuracy. Relative to statistical errors, all afterglow properties are inferred correctly to within and all kilonova properties are within . We include the complete corner plot for all inferred parameters in Fig. 11 in the appendix. These results provide additional evidence that jointly fitting a contaminated signal is the most accurate and reliable method to use for disentangling and interpreting these transients.
@@ 4.3 Scenario 3 - Data without X-ray and radio bands
We have surmised that the individual afterglow recoveries thus far benefit from the presence of X-ray and radio observations that remain free of kilonova contamination. Here, we explicitly test this by starting with the same data sets used in Sec. 4.1 and Sec. 4.2. We first remove the X-ray and radio data points, and then reattempt to recover the afterglow component with our different fitting strategies. The result of this individual afterglow fit for the on-axis system, without the extra bandpasses for calibration, predicts most parameter values to be correct to within 3. However, for many parameters the magnitude of the statistical uncertainties and therefore the width of the posterior distributions have increased. This suggests that without X-ray and radio observations to ’guide’ the inference, the precision of property predictions is reduced.
@@
The power of the uncontaminated X-ray and radio bandpasses is further emphasised when we consider the individual afterglow fit for the off-axis system as presented in Fig. 7a. This afterglow fit is much worse than that of the on-axis system, and recovers only , and correctly to within 3. When we observe the predicted lightcurve from the posteriors in Fig. 7a and compare it to the expected lightcurve based on the injection values, the obvious mismatch makes it clear why the majority of inferred properties are incorrect. As the true afterglow lightcurve shape is difficult to determine from the data, and the X-ray and radio observations are not available to provide such information, the initially kilonova dominated data points appear to be fit as an extension of the afterglow signal itself. We saw in Fig. 4 (panel E) from Sec. 4.2 that the equivalent kilonova only recovery also fits these same data points and correctly infers the transient parameters. Since the individual fits contradict each other, if confronted with real data lacking the radio and X-ray observations, it could be difficult to determine which of the individual transient recoveries are correct. This means in the case of an off-axis system, which may have contamination and is more likely missing crucial X-ray and radio observations, individual fits or those performed after subtraction are not reliable.
@@
Given the stark failure of individual fitting methods to the off-axis system, we apply the joint fitting strategy to both the on and off-axis scenarios, and compare our results. In both cases, we see improvement and find that all inferred parameters of the kilonovae and afterglows are consistent with the input values. The use of a combined model allows a correct separation of the data points into their respective transient components, which resolves the problem observed in the individual fit of Fig. 7a. However, we note that while accuracy has improved, the posterior distributions have broadened (as expected) compared to the joint fits from Secs. 4.1 and 4.2, which included X-ray and radio observations. An example of this can be seen in the appendix, where we show the full corner plot from the joint inference of the off-axis GRB system.
@@ 4.4 Scenario 4 - On-axis afterglow with little kilonova contamination
Thus far, we have shown that significant contamination between an afterglow and a kilonova can cause notable biases in parameter inference if the most common individual fitting strategies continue to be used. Here, we consider the scenario where contamination is much less significant. In this scenario, data is simulated for an on-axis GRB system with our key afterglow parameters changed to , , and , creating a particularly bright transient. The remainder of the injection values for the afterglow and kilonova remain the same as in the previous scenarios. This results in an afterglow dominated situation with very little contamination from the kilonova, which is observable as a small bump on the lightcurve at frequencies less than that of the lsst-z band.
@@
Beginning, as usual, with the individual fits, we show in Fig. 8 the violin plots for the inference of key afterglow parameters (upper pannel in red) and kilonova parameters (lower panel in blue). These plots show that there is heavy bias in both the afterglow and kilonova individual fits, despite the afterglow being the dominant component and containing X-ray and radio data for calibration. By employing Bayesian inference with the joint model, we see that all properties of both transients are consistent with the injection values. The best fit posterior lightcurve for the joint fit is shown in Fig. 9 which makes it easy to see how subtle the contamination is compared to panel C in Fig. 4. Our results emphasise that where contamination is suspected, even if not necessarily obvious, a combined model is required to allow an accurate interpretation of the transients.

### v2

@@ 4.1 Scenario 1 - On-axis afterglow with significant kilonova contamination
As discussed in Sec. 3, an on-axis GRB system is likely to suffer from significant kilonova contamination. As the cases of GRB211211A and GRB230307A already prove, they are also the most observationally relevant. Our fits using strategies 1 and 2 for this on-axis scenario and the following off-axis scenario are presented in Fig. 4, where each subplot shows 100 randomly drawn light curves from the posterior distributions as solid lines, with colours corresponding to the observation band. For comparison, we overplot the expected curves of the afterglow and kilonova components, based on their injection values, as dashed and dotted lines respectively.
@@
We first focus on the strategy where the kilonova is fit independently to the data. The fitted light curve from this strategy is shown in Fig. 4 (panel B), which highlights the poor recovery relative to the true input. However, we emphasize that in real observation scenarios, true parameter values are not known, so one could not directly assess the accuracy of inference in this manner. If we disregard the true input curves, the posteriors in Fig. 4 appear to reasonably match the relevant data points for the transient being fit, suggesting a ‘good’ fit to the data that may lead to the conclusion that the inferred parameters are correct.
@@
In Fig. 5, and similar figures within this section, we show violin plots for relevant afterglow parameters in the upper panel and kilonova parameters in the lower panel with individual afterglow fits in green, individual kilonova fits in blue, kilonova fits after afterglow subtraction in red, and joint fits in yellow. Injection values are marked by ‘’. Despite producing good visual fits to the data, our inferred parameters from the individual kilonova fit are significantly biased. In particular, , are biased by and , respectively (relative to the statistical errors), and show no posterior overlap with the true input. Given that these parameters are the most important for interpreting a kilonova, the error seen here suggests that solely fitting the kilonova is not a good strategy.
@@
As shown in Fig. 4 (panel A), the independent fit to the afterglow is a more viable strategy for producing a better match to the true input. This is largely attributed to including X-ray and radio data, which provide some “calibration” for the afterglow model as the kilonova model has no flux at these wavelengths. This motivates the use of the commonly-used strategy of attempting to isolate the kilonova signal by first subtracting the predicted afterglow, and then performing a fit. We approach this “subtraction” strategy by masking the X-ray and radio data points and then subtracting the afterglow contribution from all other wavelengths. We generate the afterglow contribution using our ground truth “tophatredback” afterglow model with the best-fitting model parameters as inputs and no additional error term. Ideally, the errors on the subtracted data points should account for the errors on both the model and the original data. The former can be estimated using samples from the posterior, and in our study for most of our simulated data, this produces errors significantly smaller than the original simulated data error. We therefore work with the original error to simplify our analyses when accounting for the uncertainty in the subtraction.
@@
While most of the posteriors using the subtraction method, shown in Fig. 5, are more consistent with the input than the independent kilonova-only fit, they are still significantly biased. Notably, again, neither the kilonova ejecta mass nor shell velocity are consistent with the input, highlighting the pitfalls of this commonly used strategy. Additionally, extinction is recovered worse and now shows no posterior overlap with the true input. We note that increasing the size of the (largely arbitrary) uncertainty on the error bars when fitting the afterglow-subtracted data could make the kilonova parameters more consistent with the input. However, from a Bayesian perspective, the “subtraction” approach is incorrect and an example of a misspecified model. In particular, when we first fit the afterglow model to data, to maximise the likelihood, the afterglow will depart from the ‘truth’, especially, but not only, when radio and X-ray data errors are high or the quality otherwise due to cadence or spectral coverage is low relative to the optical. This biased afterglow recovery is then built into the kilonova parameter estimation. Since the intrinsic kilonova and afterglow models are themselves uncertain (unlike our simulation where we know the input), the subtraction strategy can cause significant issues with interpretations. By contrast, the proper (joint inference) Bayesian approach to this problem produces kilonova and afterglow posteriors that are marginalising over the other (afterglow) model as opposed to being conditioned on it, and so they will naturally be consistent with the input (given adequate sampling). Moreover, by marginalising over the other model parameters, it is far more likely to yield robust estimated kilonova parameters, as opposed to the subtraction approach.
@@
To verify that the joint approach works, we show fits to the data in Fig. 4 (panel C) and yellow violins for comparison to the other methods in Fig. 5, highlighting consistency with the inputs. A full corner plot for the joint recovery can be found in Fig. A1 of the appendix, which confirms an accurate fit with all true kilonova and afterglow parameters recovered correctly to within 3 of the predicted values. We also emphasize that the joint-fitting approach is not significantly more computationally intensive than independently performing an afterglow-only and, subsequently, kilonova-only fit. Importantly, it does not require extra data processing to “generate” afterglow-subtracted kilonova data.
@@ 4.2 Scenario 2 - Off-axis afterglow with significant kilonova contamination
Following the same analysis approaches as Sec. 4.1, we again compare individual, afterglow subtracted, and joint fits for an off-axis GRB system. In Sec. 3 we noted that at early observation epochs, the off-axis afterglow spectral luminosity is typically low, which allows the kilonova to dominate the signal in all wavelengths. This suggests that strategies which involve individually fitting the kilonova to the data may be more suited towards the off-axis contamination scenario than the on-axis scenario. The result of our individually fit kilonova component relative to the true inputs, shown in Fig. 4 (panel E), seems to reflect this.
@@
In Fig. 6, we see the biases on ejecta mass and shell velocity are reduced in comparison to the on-axis GRB case, discussed in Sec. 4.1. However, the posterior of remains inconsistent with our input. We additionally see that opacity is incorrectly inferred, which is likely due to sensitivity to contamination around the knee of the lightcurve, as this is typically the source of constraints on opacity (alongside the peak time) for the kilonova models used in this work. The inaccuracy of the recovered posteriors, despite a seemingly good fit to relevant data points, highlights the need for caution in interpreting fits to subtracted data, and general caution for detailed parameter estimates of quantities with this fitting strategy.
@@
We now look at the individual afterglow fit to data and note that while the kilonova is largely dominant in the lower frequency bandpasses, up to 10 days in lsst-r, this is not a complete detriment to the afterglow recovery. This is again due to the inclusion of uncontaminated radio and X-ray observations which provide a level of calibration to the fit. In the upper panel of Fig. 6, we show green violin plots for the key afterglow parameters which we identified as significantly influential to contamination in Sec. 3, and additionally extinction. Most of these properties are recovered at a 3 level (relative to statistical errors). However, extinction, as seen in the uppermost violin in Fig. 6, and initial lorentz factor, have much greater biases as they are largely informed by the optical/NIR light curve. These inconsistencies again, prompt a review of methods to obtain more accurate predictions so that one may better interpret the afterglow.
@@
If we choose to subtract the best-fitting afterglow from the data to isolate the kilonova, the resulting inference not only still fails to correctly recover the input value of ejecta mass, but makes the estimation worse as can be seen by the red violins in Fig. 6. Instead, we may again turn to the use of a combined model. We show the joint fit in panel F of Fig. 4, and note that it provides a good visual fit to the true curves of both afterglow and kilonova components. The full simultaneous inference results show that every parameter is not only recovered correctly, but the predictions have also improved in accuracy. Relative to statistical errors, all afterglow properties are inferred correctly to within and all kilonova properties are within . We include the complete corner plot for all inferred parameters in Fig. A2 in the appendix. These results provide additional evidence that jointly fitting a contaminated signal is the most accurate and reliable method to use for disentangling and interpreting these transients.
@@ 4.3 Scenario 3 - Data without X-ray and radio bands
We have surmised that the individual afterglow recoveries thus far benefit from the presence of X-ray and radio observations that remain free of kilonova contamination. Here, we explicitly test this by starting with the same data sets used in Sec. 4.1 and Sec. 4.2. We first remove the X-ray and radio data points, and then reattempt to recover the afterglow component with our different fitting strategies. The result of this individual afterglow fit for the on-axis system (green violins in the lower panel of Fig. 7), without the extra bandpasses for calibration, correctly predicts most parameter values to within 3 relative to statistical errors. However, for many parameters the magnitude of the statistical uncertainties and therefore the width of the posterior distributions have increased relative to the individual inference of Sec. 4.1. This suggests that without X-ray and radio observations to ‘guide’ the inference, the precision of some posteriors are decreased.
The value of the uncontaminated X-ray and radio bandpasses is further emphasised when we consider the individual afterglow fit for the off-axis system. The afterglow only light curve fit, as presented in the left panel of Fig. 8, is far from the ground truth and only , and are recovered correctly to within 3. While the green violins showing the posteriors of select afterglow parameters in the upper panel of Fig. 7 seem to also show a successful recovery of , the statistical errors are so small that the median value of the distribution is inconsistent with the input.
@@
As the true afterglow light curve shape is difficult to determine from the data, and the X-ray and radio observations are not available to provide such information, the initially kilonova dominated data points appear to be fit as an extension of the afterglow signal itself. We saw in Fig. 4 (panel E) from Sec. 4.2 that the equivalent kilonova only recovery also fits these same data points and correctly infers the transient parameters. Since the individual fits contradict each other, if confronted with real data lacking the radio and X-ray observations, it could be difficult to determine which of the individual transient recoveries are correct. This means in the case of an off-axis system, which may have contamination and is more likely missing crucial X-ray and radio observations, individual fits or those performed after subtraction are not reliable.
@@
Given the stark failure of individual fitting methods to the off-axis system, we apply the joint fitting strategy to both the on and off-axis scenarios, and compare our results. In both cases, we see improvement and find that all inferred parameters are consistent with the input values, as demonstrated by the yellow violins in Fig. 7. The use of a combined model allows a correct separation of the data points into their respective transient components, which resolves the problem observed in the individual fit, as shown in Fig. 8. However, we note that while accuracy has improved, many of the posterior distributions have broadened (as expected) compared to the joint fits from Secs. 4.1 and 4.2, which included X-ray and radio observations. An example of this can be seen in Fig. A3 of the appendix, where we show the full corner plot from the joint inference of the off-axis GRB system.
@@ 4.4 Scenario 4 - On-axis afterglow with little kilonova contamination
Thus far, we have shown that significant contamination between an afterglow and a kilonova can cause notable biases in parameter inference if the most common individual fitting strategies continue to be used. Here, we consider the scenario where contamination is much less significant. In this scenario, data is simulated for an on-axis GRB system with our key afterglow parameters changed to , , rad and rad, creating a particularly bright transient. The remainder of the injection values for the afterglow and kilonova remain the same as in the previous scenarios. This results in an afterglow dominated situation with very little contamination from the kilonova, which is observable as a small bump on the light curve at frequencies less than that of the lsst-z band.
@@
Beginning, as usual, with the individual fits, we show in Fig. 9 the violin plots for the inference of key afterglow parameters in green and kilonova parameters in blue. These plots show that as expected, there are strong biases in the kilonova individual fits while estimates from the afterglow only fit are consistent with injection values. Given the small errors in the afterglow fit, one could again be motivated to consider the subtraction method. However, as shown in Fig. 9, this methodology fails to correctly recover the opacity. By employing Bayesian inference with the joint model, we see from the yellow violins that the inferred properties of the kilonova become consistent with the injection values and importantly, remain consistent for the afterglow. The best-fitting posterior light curve for the joint fit is shown in Fig. 10 which highlights how subtle the contamination is compared to panel C in Fig. 4. Our results emphasise that where contamination is truly present, even if not obvious, a combined model is required to allow an accurate interpretation of the transients.
4.5 Quantifying the robustness of fits assuming true contamination
Situational differences between scenarios in this paper, like viewing axis and data availability, are more easily comparable because we only minimally vary the injection parameters of the afterglow. In this subsection we quantify the robustness of joint fitting across a wider variety of potential scenarios than have been investigated so far. Importantly, we also account for different data realisations of the same scenario that could impact the posterior distributions. To accomplish this, we have selected six new sets of parameters that probe a larger range of the kilonova and afterglow prior space described in Tabs. 2 and 3, and simulated 10 data realisations for each set of parameters. This totals 60 different synthetic observations, each of which we perform all fitting strategies on: afterglow only, kilonova only, kilonova only after afterglow subtraction, and joint.
The parameters for each data set are still selected with the requirement that the afterglow and kilonova contaminate, and we intentionally use three scenarios where the afterglow is viewed on-axis and three where the afterglow is viewed off-axis. To compare the biases between each fitting method, we perform the following analysis on each of the six sets of data. For every recovery method, we first average over the posterior distributions from the 10 different data realisations which allows us to calculate the standard deviation of the average posterior distributions. We then calculate the absolute difference between the average of the posteriors for each parameter and the true injection values. The absolute difference is expressed in units of . When these analysis steps are completed for each recovery method, we calculate the difference between for the joint method and all other methods such that . Using this metric, a more negative (positive) indicates the parameter values inferred using the specified method of recovery are worse (better) than the joint recovery method. The results of this analysis are presented in Fig. 11 where the subscripts ‘kn’, ‘kn-sub’ and ‘ag’ refer to the kilonova only, kilonova only after afterglow subtraction, and afterglow only fitting methods, respectively. Different colours indicate different sets of data where sets 1-3 are contamination scenarios of on-axis afterglows and sets 4-6 are contamination scenarios of off-axis afterglows.
A violin perfectly centered about would indicate that the joint fitting method infers parameter values to the same degree of accuracy as the individual fitting methods. Compared to the kilonova fits, the afterglow parameter violins across both on and off-axis cases are generally centered closer to 0, suggesting that they can give comparable results to the joint fit, as we determined was a consequence of including X-ray and radio data. However, there are some notable outliers, especially in set 6, which indicate that the afterglow only fit can be up to 10 standard deviations worse than the joint fit at predicting the values of certain parameters, even with the inclusion of data from these crucial bandpasses. We also note that set 6 shows significant outliers in kilonova parameter space compared to the other off-axis contamination scenarios, with statistical inaccuracies from the kilonova only fit fairing more than 30 worse than the joint fit counterpart at the extreme.
As previously discussed, the kilonova only recoveries are expected to perform better in off-axis contamination scenarios than on-axis ones because the kilonova is dominant at early epochs. While Fig. 11 shows that this is generally true, with the values from sets 1-3 tending to be more negative than sets 4 and 5 on average, the severe outliers in set 6 demonstrate that an off-axis afterglow is an insufficient condition to make either the individual or afterglow subtracted methods for kilonova inference trustworthy options. Furthermore, we note that while the afterglow subtracted kilonova fits generally show more positive values compared to kilonova fits without prior data subtraction, we see an exception to this in set 3 where the values for and become more negative. This shows that using the subtraction technique does not guarantee an improvement in accuracy relative to purely individual fitting techniques.
4.6 Scenario 5 - Model misspecification with weakly-informative data and no kilonova
In this final scenario, we test the limits of our methodology and simulate an afterglow observation with no contamination i.e., there is no kilonova. This scenario is particularly relevant to claims of kilonova detections when understanding the distinction between long and short GRBs. The presence of a kilonova in the previous scenarios was well constrained given high time sampling in each bandpass, so we now reduce the number of data points to 20 and limit our observations to radio, X-ray, and LSST irgu bands. We also raise the level of noise in the observations to mimic a lower signal-to-noise observation. We simulate the afterglow with , , rad and rad. In our analysis, we still attempt all types of fit as previously to see if a kilonova can be predicted when not actually present.
Fig. 12 shows that since there is no longer contamination from a kilonova, the individual afterglow fit provides accurate results, inferring parameter values within 1 of the injection values. The posterior light curve for this fit is shown in the left panel of Fig. 13 which matches well to both the data and the ground truth curves. In this low quality data scenario, we find that the parameters relating solely to the afterglow are also recovered using the joint fit as reflected by the yellow violins in Fig. 12. However, extinction, being a shared parameter between the afterglow and kilonova in the joint fit, is biased by the attempt to recover the latter non-existent transient and shows no posterior overlap with the true input. More generally, and especially when the data is informative (more bandpasses and higher time sampling), the independent afterglow fit is the only method that can reliably recover the ground truth. This is again a product of model misspecification, i.e., a bias induced in Bayesian inference when models used in fitting are not representative of the true data generation process.
When we remove the best-fitting afterglow from the data, there are very few points remaining to perform the subtracted kilonova fit on. The resulting inference, shown by the red violins in Fig. 12, produces posteriors for all parameters that are heavily skewed to the edges of our prior space and have small or no associated statistical error, as the posteriors are strongly railing against the prior. Simply, in such a scenario there are not enough remaining points to perform a good fit and, as one might expect, the posterior light curves do not match the data well.
We find that both the individual kilonova fit and the joint fit predict a kilonova, despite that it is not present in the data, but whether the fit is convincing is highly dependent on the quality of the data. In this scenario, the data is particularly uninformative in the redder filters and between approximately 0.4 – 3 days, which aligns with the optimal kilonova observation regime that we highlighted in Sec. 3. The resulting light curve fits shown in Fig. 13 from both the individual (middle panel) and joint (right panel) fits seem sensible for this scenario, with respect to the data. However, with higher quality data in the aforementioned regime, fits can again lead to posteriors that strongly rail against the prior and posterior light curves that would raise suspicion – indicators of a misspecified model. We note that in cases like the example we have shown, where data is uninformative, the prediction of a kilonova using the joint fit may be a concern for interpreting results. In these situations, posterior predictive checks could help to determine the misspecified model. We do not detail this further here, but instead refer the reader to e.g. Gelman et al. (1996) and references therin.